A large supermarket carries four qualities of ground beef. Customers are believed to purchase these four varieties with probabilities of 0.15, 0.19, 0.11, and 0.55, respectively, from the least to most expensive variety. A sample of 515purchases resulted in sales of 47, 148, 70, and 250 of the respective qualities. Does this sample contradict the expected proportions? Use α = .05.

(a) Find the test statistic. (Give your answer correct to two decimal places.)

(ii) Find the p-value. (Give your answer bounds exactly.)
< p <  
Answer #1

applying chi square goodness of fit test:

           relative observed Expected residual Chi square
category frequency Oi Ei=total*p R2i=(Oi-Ei)/√Ei R2i=(Oi-Ei)2/Ei
1 0.150 47 77.25 -3.44 11.845
2 0.190 148 97.85 5.07 25.703
3 0.110 70 56.65 1.77 3.146
4 0.550 250 283.25 -1.98 3.903
total 1.000 515 515 44.597

from above:

a) test statistic X2 =44.60

b)

p value bound: 0.0000 < p value <0.001